# Cohomology of modular form connections on complex curves

### v3

@@ 3.9. The genus $g$ generalizations of elliptic functions
where $\mathcal{E}_{m}^{n}(y)=\sum_{\ell=0}^{2p-2}\partial^{(m)}f_{\ell}(y)\;\partial^{(n)}y^{\ell}$, $\psi_{p}^{(0)}(x,y)=\frac{1}{x-y}+\sum_{\ell=0}^{2p-2}f_{\ell}(x)y^{\ell}$, for any Laurent series $f_{\ell}(x)$ for $\ell=0$, $\ldots$, $22p-$ 2. Define the doubly indexed matrix $\Delta=(\Delta_{ab}(m,n))$ by $\Delta_{ab}(m,n)=\delta_{m,n+2p-1}\delta_{ab}$. Denote by $\widetilde{R}=R\Delta$, and the formal inverse $(I-\widetilde{R})^{-1}$ is given by $\left(I-\widetilde{R}\right)^{-1}=\sum_{k\geq 0}\widetilde{R}^{\,k}$. Define $\chi(x)=(\chi_{a}(x;\ell))$ and $o(\bm{y}_{k},\mu_{0})=(o_{a}(\bm{y}_{k};\mu_{0},\ell))$, are finite row and column vectors indexed by $a\in\mathcal{I}$, $20\leq\ell\leq 2p-$ 2 with $\chi_{a}(x;\ell)=\rho_{a}^{-\frac{\ell}{2}}(p(x)+\widetilde{p}(x)(I-\widetilde{R})^{-1}R)_{a}(\ell)$, $o_{a}(\ell)=o_{a}(\bm{y}_{k},\mu_{0},\ell)=\rho_{a}^{\frac{\ell}{2}}X_{a}(\ell)$, and where $\widetilde{p}(x)=p(x)\Delta$. $\psi_{p}(x,y)$ is defined by $\psi_{p}(x,y)=\psi_{p}^{(0)}(x,y)+\widetilde{p}(x)(I-\widetilde{R})^{-1}q(y)$. For each $a\in\mathcal{I}_{+}$ we define a vector $\theta_{a}(x)=(\theta_{a}(x;\ell))$, indexed by $20\leq\ell\leq 2p-$ 2 with components $\theta_{a}(x;\ell)=\chi_{a}(x;\ell)+(-1)^{p}\rho_{a}^{p-1-\ell}\chi_{-a}(x;2p-2-\ell)$. Now define the following vectors of formal differential forms $P(x)=p(x)\;dx^{p}$, $Q(y)=q(y)\;dy^{1-p}$, with $\widetilde{P}(x)=P(x)\Delta$. Then with
@@ 3.10. The genus $g$ Schottky case
In this Subsection we recall [45, 44] the construction and reduction relations for modular functions depending on $n$ variables defined on a genus $g$ Riemann surface $M$ formed in the Schottky parameterization. All expressions here are functions of formal variables $w_{\pm a}$, $\rho_{a}\in\mathbb{C}$. Then we recall the genus $g$ reduction formula with universal coefficients that have a geometrical meaning and are meromorphic on $M$. These coefficients are generalizations of the elliptic Weierstrass functions [28]. For a $2g$ local coordinates $\bm{w}_{2g}=(w_{-1},w_{1};\ldots;w_{-g},w_{g})$, of $2g$ points $(p_{-1},p_{1};\ldots;p_{-g},p_{g})$ on the Riemann sphere, consider the genus zero point function depending on $2g$ variables $\mathcal{Z}(\bm{w}_{2g},\mu)=\mathcal{Z}(w_{-1},w_{1};\ldots;w_{-g},w_{g},\mu)=\prod_{a\in\mathcal{I}_{+}}\rho_{a}^{\beta_{a}}\mathcal{Z}(w_{-1},w_{1};\ldots;w_{-g},w_{g},\mu)$, where +=$\mathcal{I}_{+}=\{1,2,\ldots,g\}$, and $\beta_{a}$ are certain parameters related to $\mu$. Let us denote $\bm{z}_{+}=(z_{1},\ldots,z_{g})$, $\bm{z}_{-}=(z_{-1},\ldots,z_{-g})$. Let $w_{a}$ for $a\in\mathcal{I}$ be $2g$ formal variables. One identify them with the canonical Schottky parameters (for details of the Schottky construction, see [45, 44]). One can define the genus $g$ modular function not depending on variables as $\mathcal{Z}=(\bm{w}_{2g},\bm{\rho}_{2g},\mu)=\sum_{\bm{z}_{+}}\mathcal{Z}(\bm{z}_{2g},\bm{w}_{2g},\mu)$, for $(\bm{w}_{2g},\bm{\rho}_{2g})=(w_{\pm 1},\rho_{1};\ldots;w_{\pm g},\rho_{g})$. Now we recall the formal reduction formulas for all genus $g$ Schottky functions depending on $n$ variables. One defines the genus $g$ formal modular function depending on $n$ variables for $\bm{y}_{n}$ by $\mathcal{Z}(\bm{y}_{n},\mu)=\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\bm{\rho}_{2g},\mu)=\sum_{\bm{z}_{+}}\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)$, with $\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)=\mathcal{Z}(\bm{y}_{n};\bm{w}_{-1,g},\mu)$, $\mathcal{Z}(\bm{y}_{n},\mu)=\sum_{{\bf z}_{+}\in\bm{\alpha}_{g}}\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)$, where here the sum is over a basis ${\bm{\alpha}}$. It follows that $\mathcal{Z}(\bm{y}_{n},\mu)=\sum_{\bm{\alpha}_{g}\in\bm{A}}\mathcal{Z}_{\bm{\alpha}_{g}}^{(g)}(\bm{y}_{n},\mu)$, where the sum ranges over $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{g})\in\bm{A}$, for $\bm{A}=A^{\otimes{g}}$. Finally, one defines corresponding formal differential forms depending on $n$ variables $Z(\bm{y}_{n},\mu)=\mathcal{Z}(\bm{y}_{n},\mu)\;\bm{dy^{\beta}}_{n}$, $Z_{\bm{\alpha}_{g}}(\bm{y}_{n},\mu)=\mathcal{Z}_{\bm{\alpha}}(\bm{y}_{n},\mu)\;\bm{dy^{\beta}}_{n}$, where $\bm{dy^{\beta}}_{n}=\prod_{k=1}^{n}dy_{k}^{\beta_{k}}$.
@@
In [45] is was proven that the genus $g$ formal modular differential depending on $(n+1)$ variables $Z(x;\bm{y},\mu)$, for $x_{n+1}$, and point $p_{0}$, with the coordinate $y_{n+1}$, and ${\bf p}_{n}$ with coordinates ${\bf y}_{n}$ satisfies the recursive identity for ${\bf z}_{n}=(\bm{y})$
@@
Here $\partial^{(0,j)}$ is given by $\partial^{(i,j)}f(x,y)=\partial_{x}^{(i)}\partial_{y}^{(j)}f(x,y)$, for a function $f(x,y)$, and $\partial^{(0,j)}$ denotes partial derivatives with respect to $x$ and $y_{j}$. The forms $\Psi_{p}(y_{n+1},y_{k})\;dy_{k}^{j}$ given by (3.12), where $\Theta_{a}(x)$ is of (3.13), and $O^{W_{\bm{\alpha}}}_{a}(z_{n+1},{\bf z}_{n},\mu)$ of (3.14).

### v5

@@ 3.9. The genus $g$ generalizations of elliptic functions
where $\mathcal{E}_{m}^{n}(y)=\sum_{\ell=0}^{2p-2}\partial^{(m)}f_{\ell}(y)\;\partial^{(n)}y^{\ell}$, $\psi_{p}^{(0)}(x,y)=\frac{1}{x-y}+\sum_{\ell=0}^{2p-2}f_{\ell}(x)y^{\ell}$, for any Laurent series $f_{\ell}(x)$ for $\ell=0$, $\ldots$, 2p22p-2. Define the doubly indexed matrix $\Delta=(\Delta_{ab}(m,n))$ by $\Delta_{ab}(m,n)=\delta_{m,n+2p-1}\delta_{ab}$. Denote by $\widetilde{R}=R\Delta$, and the formal inverse $(I-\widetilde{R})^{-1}$ is given by $\left(I-\widetilde{R}\right)^{-1}=\sum_{k\geq 0}\widetilde{R}^{\,k}$. Define $\chi(x)=(\chi_{a}(x;\ell))$ and $o(\bm{y}_{k},\mu_{0})=(o_{a}(\bm{y}_{k};\mu_{0},\ell))$, are finite row and column vectors indexed by $a\in\mathcal{I}$, $0\leq\ell\leq 2p-2$ with $\chi_{a}(x;\ell)=\rho_{a}^{-\frac{\ell}{2}}(p(x)+\widetilde{p}(x)(I-\widetilde{R})^{-1}R)_{a}(\ell)$, $o_{a}(\ell)=o_{a}(\bm{y}_{k},\mu_{0},\ell)=\rho_{a}^{\frac{\ell}{2}}X_{a}(\ell)$, and where $\widetilde{p}(x)=p(x)\Delta$. $\psi_{p}(x,y)$ is defined by $\psi_{p}(x,y)=\psi_{p}^{(0)}(x,y)+\widetilde{p}(x)(I-\widetilde{R})^{-1}q(y)$. For each $a\in\mathcal{I}_{+}$ we define a vector $\theta_{a}(x)=(\theta_{a}(x;\ell))$, indexed by $0\leq\ell\leq 2p-2$ with components $\theta_{a}(x;\ell)=\chi_{a}(x;\ell)+(-1)^{p}\rho_{a}^{p-1-\ell}\chi_{-a}(x;2p-2-\ell)$. Now define the following vectors of formal differential forms $P(x)=p(x)\;dx^{p}$, $Q(y)=q(y)\;dy^{1-p}$, with $\widetilde{P}(x)=P(x)\Delta$. Then with
@@ 3.10. The genus $g$ Schottky case
In this Subsection we recall [46, 45] the construction and recursion relations for correlation functions depending on $n$ variables defined on a genus $g$ Riemann surface $M$ formed in the Schottky parameterization. All expressions here are functions of formal variables $w_{\pm a}$, $\rho_{a}\in\mathbb{C}$. Then we recall the genus $g$ recursion formula with universal coefficients that have a geometric meaning and are meromorphic on $M$. These coefficients are generalizations of the elliptic Weierstrass functions [29]. For a $2g$ local coordinates $\bm{w}_{2g}=(w_{-1},w_{1};\ldots;w_{-g},w_{g})$, of $2g$ points $(p_{-1},p_{1};\ldots;p_{-g},p_{g})$ on the Riemann sphere, consider the genus zero point function depending on $2g$ variables $\mathcal{Z}(\bm{w}_{2g},\mu)=\mathcal{Z}(w_{-1},w_{1};\ldots;w_{-g},w_{g},\mu)=\prod_{a\in\mathcal{I}_{+}}\rho_{a}^{\beta_{a}}\mathcal{Z}(w_{-1},w_{1};\ldots;w_{-g},w_{g},\mu)$, where $\mathcal{I}_{+}=\{1,2,\ldots,g\}$, and $\beta_{a}$ are certain parameters related to $\mu$. Let us denote $\bm{z}_{+}=(z_{1},\ldots,z_{g})$, $\bm{z}_{-}=(z_{-1},\ldots,z_{-g})$. Let $w_{a}$ for $a\in\mathcal{I}$ be $2g$ formal variables. One identify them with the canonical Schottky parameters (for details of the Schottky construction, see [46, 45]). One can define the genus $g$ correlation function not depending on variables as $\mathcal{Z}=(\bm{w}_{2g},\bm{\rho}_{2g},\mu)=\sum_{\bm{z}_{+}}\mathcal{Z}(\bm{z}_{2g},\bm{w}_{2g},\mu)$, for $(\bm{w}_{2g},\bm{\rho}_{2g})=(w_{\pm 1},\rho_{1};\ldots;w_{\pm g},\rho_{g})$. Now we recall the formal recursion formulas for all genus $g$ Schottky functions depending on $n$ variables. One defines the genus $g$ formal correlation function depending on $n$ variables for $\bm{y}_{n}$ by $\mathcal{Z}(\bm{y}_{n},\mu)=\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\bm{\rho}_{2g},\mu)=\sum_{\bm{z}_{+}}\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)$, with $\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)=\mathcal{Z}(\bm{y}_{n};\bm{w}_{-1,g},\mu)$, $\mathcal{Z}(\bm{y}_{n},\mu)=\sum_{{\bf z}_{+}\in\bm{\alpha}_{g}}\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)$, where here the sum is over a basis $\bm{\alpha}$}. It follows that $\mathcal{Z}(\bm{y}_{n},\mu)=\sum_{\bm{\alpha}_{g}\in\bm{A}}\mathcal{Z}_{\bm{\alpha}_{g}}^{(g)}(\bm{y}_{n},\mu)$, where the sum ranges over $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{g})\in\bm{A}$, for $\bm{A}=A^{\otimes{g}}$. Finally, one defines corresponding formal differential forms depending on $n$ variables $Z(\bm{y}_{n},\mu)=\mathcal{Z}(\bm{y}_{n},\mu)\;\bm{dy^{\beta}}_{n}$, $Z_{\bm{\alpha}_{g}}(\bm{y}_{n},\mu)=\mathcal{Z}_{\bm{\alpha}}(\bm{y}_{n},\mu)\;\bm{dy^{\beta}}_{n}$, where $\bm{dy^{\beta}}_{n}=\prod_{k=1}^{n}dy_{k}^{\beta_{k}}$.
@@
In [46] is was proven that the genus $g$ formal correlation differential depending on $(n+1)$ variables $Z(x;\bm{y},\mu)$, for $x_{n+1}$, and point $p_{0}$, with the coordinate $y_{n+1}$, and ${\bf p}_{n}$ with coordinates ${\bf y}_{n}$ satisfies the recursive identity for ${\bf z}_{n}=(\bm{y})$
@@
Here $\partial^{(0,j)}$ is given by $\partial^{(i,j)}f(x,y)=\partial_{x}^{(i)}\partial_{y}^{(j)}f(x,y)$, for a function $f(x,y)$, and $\partial^{(0,j)}$ denotes partial derivatives with respect to $x$ and $y_{j}$. The forms $\Psi_{p}(y_{n+1},y_{k})\;dy_{k}^{j}$ given by (LABEL:psih), where $\Theta_{a}(x)$ is of (LABEL:thetanew), and $O^{W_{\bm{\alpha}}}_{a}(z_{n+1},{\bf z}_{n},\mu)$ of (LABEL:oat).